A particle has an initial velocity of 100 m/s up to the right at 30° with the horizontal. The components of the acceleration are constant at ax=-4m/s^2 and ay=-20m/s^2. Computing the horizontal distance covered until the particle reaches a point 60m below its original elevation would be

Grade:12th pass

1 Answers

Sahithya
4 years ago
Vo=100m/s
Vox=100cos30°=86.6m/s
Voy=100sin30°=50m/s
ay=-20m/s^2
Sy=voyt+1/2ayt^2
-60=50t-1/2(20) t^2
t=6s
ax=-4m/s^2
Sx=voxt+1/2axt^2
Sx=86.6(6) +1/2(-4) (6×6) 
Sx=448m